# Birational Geometry of Special Quotient Foliations and Chazy’s Equations

### v1

@@ 6.3. The quotient foliations
It is preserved by the standard quadratic Cremona map $Q$. There is a remarkable configuration in $\mathbb{P}^{2}$ associated to the Cremona involution, the complete quadrangle associated to its four fixed points, presented in Figure 15. In it we have, in red, the four fixed points of $Q$, placed at $(1:1:1)$, $(-1:1:1)$, $(1:-1:1)$ and $(1:1:-1)$. The dashed green lines are those of the configuration $\mathcal{L}_{6}$, the six lines joining pairs of fixed points of $Q$, along which the Jacobian of $\Pi$ (31) vanishes. They are transverse to the foliation and preserved by $Q$, which restricts to each one of them as an involution. The pairs of lines of $\mathcal{L}_{6}$ that do not share a common fixed point of $Q$ intersect, by pairs, at the three indeterminacy points of $Q$, $(1:0:0)$, $(0:0:1)$ and $(0:0:1)$. The three lines through these points, in blue in Figure 15, are those of the coordinate triangle $\Delta_{3}:XYZ=0$, and are $\mathcal{F}_{\lambda}$-invariant. The Cremona involution $Q$ contracts these lines, while blowing up the vertices of the triangle; it exchanges a line of the triangle with its opposite vertex.
@@
Figure 16 focuses in the coordinate triangle $\Delta_{3}$ and the eigenvalues of the vector field at the singular points of $\mathcal{F}_{\lambda}$.
@@
The foliation $\mathcal{F}_{\lambda}$ induces a foliation on the quotient of $\mathbb{P}^{2}$ under the action of $Q$. In the birational plane model for this quotient given by $\Pi$, this foliation is the foliation $\mathcal{G}_{\lambda}$ given by the form (12), a fact that can be established trough a direct calculation, proving Theorem 10. Let us nevertheless explain how the form (12) was obtained, and describe the geometry of both $\Pi$ and the foliation $\mathcal{G}_{\lambda}$.
Consider the vector space $\Omega_{3}$ of one-forms on $\mathbb{C}^{3}$ of the form $A\,\mathrm{d}x+B\,\mathrm{d}y+C\,\mathrm{d}z$, with $A$, $B$ and $C$ homogeneous polynomials of degree four in $x,y,z$. Consider also the linear homogeneous vector fields $E=X\,\partial/\partial X+Y\,\partial/\partial Y+Z\,\partial/\partial Z$ and $V=X\,\partial/\partial X+\lambda^{2}Y\,\partial/\partial Y-\lambda Z\,\partial/\partial Z$, which are linearly independent on a Zariski-open subset, and which are in the kernel of the form (32) generating $\mathcal{F}_{\lambda}$. Within $\Omega_{3}$, the elements $w$ for which the conditions $(\Pi^{*}w)E\equiv 0$ and $(\Pi^{*}w)V\equiv 0$ hold is a linear subspace, that may be defined by explicit linear equations on the coefficients of the polynomials $A$, $B$ and $C$. By solving this system, this subspace is found to have dimension one, and to be generated by the form (12).
@@
$xyz(x+y+z)=0,$
(in red in Figure 17), together with its six vertices, the points of intersection of each pair of lines of the quadrilateral: $(0:0:1)$, $(0:1:0)$, $(1:0:0)$, $(-1:0:1)$, $(0:-1:1)$ and $(-1:1:0)$, in green in Figure 17. These six points come in three pairs (pairs without a common line), and each pair determines a line; these are the diagonals
@@
in blue in Figure 17. The three points of intersection of the diagonals are the diagonal points, $(-1:1:1)$, $(1:-1:1)$ and $(1:1:-1)$. Each line of the quadrilateral is incident to three vertices. Each diagonal is incident to two vertices and two diagonal points. Each vertex lies on two sides of the quadrilateral and one diagonal. Each diagonal point lies on two diagonals.
@@
The seven lines of this configuration are tangent to the foliation $\mathcal{G}_{\lambda}$, and its nine points are singular points of the foliation, which has four further singular points, one on each line of the quadrilateral, with eigenvalues $-2:1$, and whose position depends upon $\lambda$. These account for all thirteen singular points of the degree three foliation $\mathcal{G}_{\lambda}$, the number of singularities of a degree three foliation in the plane (counted with multiplicity).
@@
By blowing up the six vertices of the quadrilateral (the triple points of the configuration) each one of its four lines becomes a curve of self-intersection $-2$, corresponding to the resolution of a singularity of type $A_{1}$. The triangle formed by the diagonals becomes a cycle of three curves of self-intersection $-1$.
@@ 7. The groups of birational automorphisms of the special quotient foliations
Theorem 5 will be proved in this section. To a holomorphic foliation by curves on an algebraic manifold, we can associate the space of its Zariski-dense leaves, a (not-necessarily-Hausdorff) complex manifold. This space is a birational invariant, and the group of birational transformations preserving the foliation acts holomorphically on it. In the cases we consider, these spaces are either elliptic curves or their quotients under the action of a finite group acting with fixed points (elliptic orbifolds), and, for all of these, the groups of biholomorphisms can be easily described. This will be the starting point for Theorem 5.
@@ 7.1. Birational symmetries of linear foliations
Let =$\mathbb{H}=\{\tau\in\mathbb{C}\mid\Im(\tau)>0\}$. For $\tau\in\mathbb{H}$, let $\mathcal{L}_{\tau}$ be the linear foliation on $\mathbb{P}^{2}$ given in the chart $(x:y:1)$ by the vector field $X_{\tau}=\tau x\,\partial/\partial x+y\,\partial/\partial y$. Consider two actions of $\mathrm{SL}(2,\mathbb{Z})$: the first, by fractional linear transformations on $\mathbb{H}$,
@@
With respect to these, for $A\in\mathrm{SL}(2,\mathbb{Z})$,
$A_{*}\mathcal{L}_{\tau}=\mathcal{L}_{A\cdot\tau}.$
@@
In this setting, we may describe the group of birational automorphisms of $\mathbb{P}^{2}$ that preserve $\mathcal{L}_{\tau}$:
@@
Let $\tau\in\mathbb{H}$. Let $\mathrm{Bir}(\mathbb{P}^{2},\mathcal{L}_{\tau})$ be the group of birational transformations of $\mathbb{P}^{2}$ that preserve $\mathcal{L}_{\tau}$. Let $G_{\tau}^{0}\subset\mathrm{PGL}(3,\mathbb{C})\cap\mathrm{Bir}(\mathbb{P}^{2},\mathcal{L}_{\tau})$ be the subgroup generated by the flows of $X_{\tau}$ and $x\,\partial/\partial x$. Let $I_{\tau}\subset\mathrm{SL}(2,\mathbb{Z})$ be the stabilizer of $\tau$ under the action (33). The group $\mathrm{Bir}(\mathbb{P}^{2},\mathcal{L}_{\tau})$ is the semidirect product $I_{\tau}\ltimes G_{\tau}^{0}$, with $I_{\tau}$ acting birationally on $\mathbb{P}^{2}$ via the monomial action (34).
The result will be straightforward consequence of the upcoming Proposition 19 and its proof. The spaces of transcendental leaves will establish a link between the problem at hand and that of the classification of elliptic curves.
Let $\Lambda_{\tau}\subset\mathbb{C}$ the lattice generated by $1$ and $\tau$, and let $E_{\tau}$ be the elliptic curve $\mathbb{C}/\Lambda_{\tau}$. The group of deck transformations of the universal covering $\mathbb{C}\to E_{\tau}$ is isomorphic to $\mathbb{Z}^{2}$, and given by
@@
which maps the vector field $\partial/\partial w$ to $-2\mathrm{i}\pi X_{\tau}$, and for which
@@
which act on $w$ by translations. The projection $z$ is equivariant with respect to the action (35) and to this last one: $\gamma_{m,n}\circ z=z\circ\overline{\gamma}_{m,n}$. This establishes the claim. ∎
@@
The action of $\mathrm{Bir}(\mathbb{P}^{2},\mathcal{L}_{\tau})$ on the space of transcendental leaves of $\mathcal{L}_{\tau}$ will give the homomorphism from $\mathrm{Bir}(\mathbb{P}^{2},\mathcal{L}_{\tau})$ to $\mathrm{Bih}(E_{\tau})$ associated to the decomposition (36). We begin by showing that every element of $\mathrm{Bir}(\mathbb{P}^{2},\mathcal{L}_{\tau})$ is holomorphic in restriction to $U$. Let $L\subset U$ be a leaf of $\mathcal{L}_{\tau}$. It is an entire, Zariski-dense curve, parametrized by $\mathbb{C}$ as a solution to $X_{\tau}$; in particular, it has a natural affine coordinate. Let $\sigma\in\mathrm{Bir}(\mathbb{P}^{2},\mathcal{L}_{\tau})$, and let $\Omega_{\sigma}\subset\mathbb{P}^{2}$ be a Zariski-open subset in restriction to which $\sigma$ is a biholomorphism onto its image. The restriction of $\sigma$ to $L\cap\Omega_{\sigma}$ extends as a holomorphic map from $L$ to $\mathbb{P}^{2}$, and its image is an entire transcendental curve tangent to $\mathcal{L}_{\rho}$, another one of its leaves, contained in $U$. In restriction to $L$, and with respect to the global affine coordinates both in $L$ and in the curve into which $L$ is mapped, $\sigma$ is an affine map. Let $\Phi:\mathbb{C}\times U\to U$ denote the restriction to $U$ of the flow of $X_{\tau}$. Let $D\subset\mathbb{C}$ be the unit disk, and $j:D\to\Omega_{\sigma}$ a sufficiently small transversal to $\mathcal{L}_{\rho}$ intersecting $L$. We have a covering tube $\psi:D\times\mathbb{C}\to U$, $\psi(z,t)=\Phi(t,j(z))$, that glues holomorphically the leaves of $\mathcal{L}_{\rho}$ intersecting the transversal (in general, such tubes glue universal coverings of leaves; they have been considered by Brunella in connection with the problem of simultaneous uniformization [Bru11]). We have that $\psi$ is a biholomorphism onto its image. The covering tube exhibits the fact that the affine structures along the leaves of $\mathcal{L}_{\rho}$ vary holomorphically in the direction transverse to the foliation (that we have a foliated affine structure in the sense of [DG23]). By the previous discussion, there exist holomorphic functions $\alpha$ and $\beta$, with $\alpha$ non-vanishing, such that for every $(z,t)\in D\times\mathbb{C}$,
@@
$0\to\mathbb{C}\to\mathrm{Bih}(U,\mathcal{L}_{\tau}|_{U})\to\mathrm{Bih}(E_{\tau})\to 0,$
in which the subgroup of $\mathrm{Bih}(U,\mathcal{L}_{\tau}|_{U})$ of transformations that induce trivial biholomorphisms of $E_{\tau}$ is given by the flow of $X_{\tau}$. The vector field $x\,\partial/\partial x$ preserves $U$ and $\mathcal{L}_{\tau}$, and induces, via $f_{\tau}$, a non-trivial holomorphic vector field on $E_{\tau}$, which generates its group of translations. We have the action of $I_{\tau}$ on $E_{\tau}$ induced, for $A=\left(\begin{array}[]{cc}a&b\\ c&d\end{array}\right)\in I_{\tau}$, by multiplication by $(c\tau+d)^{-1}$ on $\mathbb{C}$. With respect to the action of $I_{\tau}$ on $U$ given by the restriction of the monomial action (34), $f_{\tau}$ is equivariant:
@@
with $\alpha$ and $\beta$ holomorphic functions ($\alpha$ a nowhere-vanishing one). In order for such a $\widetilde{\sigma}$ to induce a biholomorphism of $U$, for every $v\in\mathbb{Z}^{2}$ there must exist $v^{\prime}\in\mathbb{Z}^{2}$ such that $\widetilde{\sigma}\circ\overline{\gamma}_{v}$ and $\overline{\gamma}_{v^{\prime}}\circ\widetilde{\sigma}$ coincide. The actions on $\widetilde{E}_{\tau}$ of these are those of $\overline{\gamma}_{v}$ and $\overline{\gamma}_{v^{\prime}}$, and we should have that $v=v^{\prime}$, this is, $\widetilde{\sigma}$ should commute with all the deck transformations. This commutativity is equivalent to the condition that, for all $n,m\in\mathbb{Z}$,
(39) $\alpha(z)$ $=\alpha(z+m+n\tau),$
@@
This implies that both $\alpha$ and $\beta^{\prime}$ are holomorphic elliptic functions with periods in $\Lambda_{\tau}$, that they are both constant. If $\beta(z)=az+b$, with $a,b\in\mathbb{C}$, condition (40) is equivalent to the fact that, for all $(m,n)\in\mathbb{Z}^{2}$, $(1-\alpha-a\tau)m=an$, imposing the conditions $a=0$ and $\alpha\equiv 1$. The resulting transformations, of the form $(z,w)\mapsto(z,w+b)$, map under $\pi_{\tau}$ to transformations in the flow of $X_{\tau}$. This establishes the proposition, as well as Theorem 17. ∎

### v2

@@ 6.2. The quotients of degree-one foliations
It is preserved by the standard quadratic Cremona map $Q$. There is a remarkable configuration in $\mathbb{P}^{2}$ associated to the Cremona involution, the complete quadrangle associated to its four fixed points, presented in Figure 15. In it we have, in red, the four fixed points of $Q$, placed at $(1:1:1)$, (1:1:1)(-1:1:1), (1:1:1)(1:-1:1) and (1:1:1)(1:1:-1). The dashed green lines are those of the configuration $\mathcal{L}_{6}$, the six lines joining pairs of fixed points of $Q$, along which the Jacobian (33) of $\Pi$ vanishes. They are transverse to the foliation and preserved by $Q$, which restricts to each one of them as an involution. The pairs of lines of $\mathcal{L}_{6}$ that do not share a common fixed point of $Q$ intersect, by pairs, at the three indeterminacy points of $Q$, $(1:0:0)$, $(0:0:1)$ and $(0:0:1)$, the points in blue in Figure 15. The three lines through these points, also in blue, are those of the coordinate triangle $\Delta_{3}:XYZ=0$, and are $\mathcal{F}_{\lambda}$-invariant. The Cremona involution $Q$ contracts these lines, while blowing up the vertices of the triangle; it exchanges a line of the triangle with its opposite vertex. Figure 16 focuses in the coordinate triangle $\Delta_{3}$ and the eigenvalues of the vector field at the singular points of $\mathcal{F}_{\lambda}$.
The foliation $\mathcal{F}_{\lambda}$ induces a foliation on the quotient of $\mathbb{P}^{2}$ under the action of $Q$. In the birational plane model for this quotient given by the map $\Pi$ (Proposition 19), this foliation is the foliation $\mathcal{G}_{\lambda}$ given by the form (12), a fact that can be established through a direct calculation: the pull-back of the form (12) via the map (32) is the form (34). ∎
@@
Let us explain how the one-form (12) was obtained, and describe the geometry of both $\Pi$ and of the foliation $\mathcal{G}_{\lambda}$. Consider the vector space $\Omega_{3}$ of one-forms on $\mathbb{C}^{3}$ of the form $A\,\mathrm{d}x+B\,\mathrm{d}y+C\,\mathrm{d}z$, with $A$, $B$ and $C$ homogeneous polynomials of degree four in $x$, $y$, and $z$. Consider also the linear homogeneous vector fields
$E=X\,\frac{\partial}{\partial X}+Y\,\frac{\partial}{\partial Y}+Z\,\frac{\partial}{\partial Z}\text{ and }V=X\,\frac{\partial}{\partial X}+\lambda^{2}Y\,\frac{\partial}{\partial Y}-\lambda Z\,\frac{\partial}{\partial Z},$
which are linearly independent on a Zariski-open subset, and which are in the kernel of the form (34) generating $\mathcal{F}_{\lambda}$. Within $\Omega_{3}$, the elements $w$ for which the conditions $(\Pi^{*}w)E\equiv 0$ and $(\Pi^{*}w)V\equiv 0$ hold form a linear subspace, that may be defined by explicit linear equations on the coefficients of the polynomials $A$, $B$ and $C$. By solving this system, this subspace is found to have dimension one, and to be generated by the form (12).
@@
$xyz(x+y+z)=0$
(in red in Figure 17), together with its six vertices, the points of intersection of each pair of lines of the quadrilateral: $(0:0:1)$, $(0:1:0)$, $(1:0:0)$, (1:0:1)(-1:0:1), (0:1:1)(0:-1:1) and (1:1:0)(-1:1:0), in green in Figure 17. These six points come in three pairs (pairs without a common line), and each pair determines a line; these are the diagonals
@@
in blue in Figure 17. The three points of intersection of the diagonals are the diagonal points, (1:1:1)(-1:1:1), (1:1:1)(1:-1:1) and (1:1:1)(1:1:-1). Each line of the quadrilateral is incident to three vertices. Each diagonal is incident to two vertices and two diagonal points. Each vertex lies on two sides of the quadrilateral and one diagonal. Each diagonal point lies on two diagonals.
@@
The seven lines of this configuration are tangent to the foliation $\mathcal{G}_{\lambda}$, and its nine points are singular points of the foliation, which has four further singular points, one on each line of the quadrilateral, with eigenvalues 2:1-2:1, and whose position depends upon $\lambda$. These account for all thirteen singular points of the degree three foliation $\mathcal{G}_{\lambda}$, the number of singularities of a degree three foliation in the plane (counted with multiplicity).
@@
By blowing up the six vertices of the quadrilateral (the triple points of the configuration) each one of its four lines becomes a curve of self-intersection 2-2, corresponding to the resolution of a singularity of type $A_{1}$. The triangle formed by the diagonals becomes a cycle of three curves of self-intersection 1-1.
@@ 7. The groups of birational automorphisms of the special quotient foliations
Theorem 5 will be proved in this section. To a holomorphic foliation by curves on an algebraic surface, we can associate its leaf space, the space resulting from identifying transversals to the foliation by the holonomy relation. It is a complex not-necessarily-Hausdorff manifold. If not every leaf of the foliation is contained in an algebraic curve, then, by a theorem of Jouanolou and Ghys [Ghy00], there are at most finitely many algebraic curves invariant by the foliation. The leaves that are not contained in an algebraic curve (the Zariski-dense leaves) form an open subset of the leaf space. This space of Zariski-dense leaves is a birational invariant of the foliated surface, and the group of birational transformations preserving the foliation acts holomorphically on it. In the cases we consider, these spaces are either elliptic curves or their quotients under the action of a finite group acting with fixed points (elliptic orbifolds), and, for all of these, the groups of biholomorphisms can be easily described. This will be the starting point for the proof of Theorem 5.
@@ 7.1. Birational symmetries of linear foliations
Let $\mathbb{H}=\{\tau\in\mathbb{C}\mid\Im(\tau)>0\}$. For $\tau\in\mathbb{H}$, let $\mathcal{L}_{\tau}$ be the linear foliation on $\mathbb{P}^{2}$ given in the chart $(x:y:1)$ by the vector field $X_{\tau}=\tau x\,\partial/\partial x+y\,\partial/\partial y$. Consider two actions of $\mathrm{SL}(2,\mathbb{Z})$: the first, by fractional linear transformations on $\mathbb{H}$,
@@
With respect to these, for $A\in\mathrm{SL}(2,\mathbb{Z})$, $A_{*}X_{\tau}=(c\tau+d)X_{A\cdot\tau}$, and thus
(37) $A_{*}\mathcal{L}_{\tau}=\mathcal{L}_{A\cdot\tau}.$
@@
We may describe the group of birational automorphisms of $\mathbb{P}^{2}$ that preserve $\mathcal{L}_{\tau}$:
@@
Let $\tau\in\mathbb{H}$. Let $\mathrm{Bir}(\mathbb{P}^{2},\mathcal{L}_{\tau})$ be the group of birational transformations of $\mathbb{P}^{2}$ that preserve $\mathcal{L}_{\tau}$. Let $G_{0}^{\tau}\subset\mathrm{PGL}(3,\mathbb{C})\cap\mathrm{Bir}(\mathbb{P}^{2},\mathcal{L}_{\tau})$ be the subgroup generated by the flows of $X_{\tau}$ and $x\,\partial/\partial x$. Let $I_{\tau}\subset\mathrm{SL}(2,\mathbb{Z})$ be the stabilizer of $\tau$ under the action (35). The group $\mathrm{Bir}(\mathbb{P}^{2},\mathcal{L}_{\tau})$ is the semidirect product $I_{\tau}\ltimes G_{0}^{\tau}$, with $I_{\tau}$ acting birationally on $\mathbb{P}^{2}$ via the monomial action (36).
The result will be a consequence of the upcoming Proposition 24 and its proof. The spaces of Zariski-dense leaves will establish a link between the problem at hand and that of the classification of elliptic curves and their biholomorphisms.
Let $\Lambda_{\tau}\subset\mathbb{C}$ the lattice generated by $1$ and $\tau$, and let $E_{\tau}$ be the elliptic curve $\mathbb{C}/\Lambda_{\tau}$. The group of deck transformations of the universal covering $\mathbb{C}\to E_{\tau}$ is isomorphic to $\mathbb{Z}^{2}$, and formed by the transformations
@@
which maps the vector field $\partial/\partial w$ to $2\mathrm{i}\pi X_{\tau}$, and for which
@@
which act on $w$ by translations. The projection $\rho$ onto the first factor is equivariant with respect to the action (38) and to this last one: $\gamma_{m,n}\circ\rho=\rho\circ\overline{\gamma}_{m,n}$. This establishes the claim. ∎
Proposition 23.
For $\tau,\tau^{\prime}\in\mathbb{H}$, the foliations $\mathcal{L}_{\tau}$ and $\mathcal{L}_{\tau^{\prime}}$ are birationally equivalent if and only if $\tau$ and $\tau^{\prime}$ are in the same orbit of the action of $\mathrm{SL}(2,\mathbb{Z})$ on $\mathbb{H}$.
Proof.
If $\tau$ and $\tau^{\prime}$ are in the same orbit of the action of $\mathrm{SL}(2,\mathbb{Z})$ on $\mathbb{H}$, by (37), $\mathcal{L}_{\tau}$ and $\mathcal{L}_{\tau^{\prime}}$ are birationally equivalent. If $\mathcal{L}_{\tau}$ and $\mathcal{L}_{\tau^{\prime}}$ are birationally equivalent, their spaces of Zariski-dense leaves are biholomorphic. By the previous claim, these leaf spaces are $E_{\tau}$ and $E_{\tau^{\prime}}$, which are biholomorphic if and only if $\tau$ and $\tau^{\prime}$ are in the same orbit of the action of $\mathrm{SL}(2,\mathbb{Z})$ on $\mathbb{H}$. ∎
@@
The action of $\mathrm{Bir}(\mathbb{P}^{2},\mathcal{L}_{\tau})$ on the space of Zariski-dense leaves of $\mathcal{L}_{\tau}$ will give the homomorphism from $\mathrm{Bir}(\mathbb{P}^{2},\mathcal{L}_{\tau})$ to $\mathrm{Bih}(E_{\tau})$ associated to the decomposition (39).
We begin by showing that every element of $\mathrm{Bir}(\mathbb{P}^{2},\mathcal{L}_{\tau})$ is holomorphic in restriction to $U$. Let $L\subset U$ be a leaf of $\mathcal{L}_{\tau}$. It is an entire, Zariski-dense curve, parametrized by $\mathbb{C}$ as a solution to $X_{\tau}$; in particular, it has a natural affine coordinate. Let $\sigma\in\mathrm{Bir}(\mathbb{P}^{2},\mathcal{L}_{\tau})$, and let $\Omega_{\sigma}\subset\mathbb{P}^{2}$ be a Zariski-open subset in restriction to which $\sigma$ is a biholomorphism onto its image. The restriction of $\sigma$ to $L\cap\Omega_{\sigma}$ extends as a holomorphic map from $L$ to $\mathbb{P}^{2}$, and its image is an entire transcendental curve tangent to $\mathcal{L}_{\rho}$, another one of its leaves, contained in $U$. In restriction to $L$, and with respect to the global affine coordinates both in $L$ and in the curve into which $L$ is mapped, $\sigma$ is an affine map. Let $\Phi:\mathbb{C}\times U\to U$ denote the restriction to $U$ of the flow of $X_{\tau}$. Let $D\subset\mathbb{C}$ be the unit disk, and $j:D\to\Omega_{\sigma}$ a sufficiently small transversal to $\mathcal{L}_{\rho}$ intersecting $L$. We have a covering tube $\psi:D\times\mathbb{C}\to U$, $\psi(z,t)=\Phi(t,j(z))$, that glues holomorphically the leaves of $\mathcal{L}_{\rho}$ intersecting the transversal; [^2] Such tubes have been considered by Ilyashenko and Brunella in connection with the problem of the simultaneous uniformization of the leaves of a foliation (see [Ily98], [Bru11]). we have that $\psi$ is a biholomorphism onto its image. The covering tube exhibits the fact that the affine structures along the leaves of $\mathcal{L}_{\rho}$ vary holomorphically in the direction transverse to the foliation (that we have a foliated affine structure in the sense of [DG23]). By the previous discussion concerning the effect of $\sigma$ on a single leaf, there exist holomorphic functions $\alpha$ and $\beta$ on $D$, with $\alpha$ non-vanishing, such that for every $(z,t)\in D\times\mathbb{C}$,
@@
(41) $0\to\mathbb{C}\to\mathrm{Bih}(U,\mathcal{L}_{\tau}|_{U})\to\mathrm{Bih}(E_{\tau})\to 0,$
in which the subgroup of $\mathrm{Bih}(U,\mathcal{L}_{\tau}|_{U})$ of transformations that induce trivial biholomorphisms of $E_{\tau}$ is given by the flow of $X_{\tau}$. The vector field $x\,\partial/\partial x$ preserves $U$ and $\mathcal{L}_{\tau}$, and induces, via $f_{\tau}$, the holomorphic vector field $\partial/\partial z$ on $E_{\tau}$, which generates its group of translations. We have the action of $I_{\tau}$ on $E_{\tau}$ induced, for $A=\left(\begin{array}[]{cc}a&b\\ c&d\end{array}\right)\in I_{\tau}$, by multiplication by $(c\tau+d)^{-1}$ on $\mathbb{C}$. With respect to the action of $I_{\tau}$ on $U$ given by the restriction of the monomial action (36), $f_{\tau}$ is equivariant:
@@
with $\alpha$ and $\beta$ holomorphic functions ($\alpha$ a nowhere-vanishing one). In order for such a $\widetilde{\sigma}$ to induce a biholomorphism of $U$, for every $v\in\mathbb{Z}^{2}$ there must exist $v^{\prime}\in\mathbb{Z}^{2}$ such that $\widetilde{\sigma}\circ\overline{\gamma}_{v}$ and $\overline{\gamma}_{v^{\prime}}\circ\widetilde{\sigma}$ coincide. The actions on $\widetilde{E}_{\tau}$ of $\widetilde{\sigma}\circ\overline{\gamma}_{v}$ and $\overline{\gamma}_{v^{\prime}}\circ\widetilde{\sigma}$ match those of $\overline{\gamma}_{v}$ and $\overline{\gamma}_{v^{\prime}}$, and we should have that $v=v^{\prime}$, this is, $\widetilde{\sigma}$ should commute with all the deck transformations. This commutativity is equivalent to the condition that, for all $n,m\in\mathbb{Z}$,
$\alpha(z)$ $=\alpha(z+m+n\tau),$
@@
This implies that both $\alpha$ and $\beta^{\prime}$ are holomorphic elliptic functions with periods in $\Lambda_{\tau}$, that they are both constant. If $\beta(z)=az+b$, with $a,b\in\mathbb{C}$, condition (43) is equivalent to the fact that, for all $(m,n)\in\mathbb{Z}^{2}$, $(1-\alpha-a\tau)m=an$, imposing the conditions $a=0$ and $\alpha\equiv 1$. The resulting transformations, of the form $(z,w)\mapsto(z,w+b)$, map under $\pi_{\tau}$ to transformations in the flow of $X_{\tau}$. This establishes the proposition. ∎
@@
Proof of Theorem 21.
Let $\tau\in\mathbb{H}$. Let $\mu_{\tau}=\{\alpha\in\mathbb{C}^{*}\mid\alpha\Lambda_{\tau}=\Lambda_{\tau}\}$. It is a subgroup of $\mathbb{C}^{*}$ that acts naturally on $E_{\tau}$. The group $\mathrm{Bih}(E_{\tau})$ of biholomorphisms of $E_{\tau}$ is the semidirect product $\mu_{\tau}\ltimes E_{\tau}$, where $E_{\tau}$ acts on itself by translations [BHPV04, Ch. 5, §5]. Within $\mathrm{Bir}(\mathbb{P}^{2},\mathcal{L}_{\tau})$, $G_{0}^{\tau}$ is normalized by $I_{\tau}$. Since the factor $\mathbb{C}$ in the short exact sequence (41) belongs to $G_{0}^{\tau}$, then, in order to prove the theorem, it is sufficient to prove that every element of $\mathrm{Bih}(E_{\tau})$ is induced by an element in the subgroup of $\mathrm{Bir}(\mathbb{P}^{2},\mathcal{L}_{\tau})$ generated by $G_{0}^{\tau}$ and $I_{\tau}$. As we discussed in the previous proof, we have an onto map $G_{0}^{\tau}\to E_{\tau}$ (taking values in the group pf translations of $E_{\tau}$), and an isomorphism between $I_{\tau}$ and $\mu_{\tau}$, both induced by the respective actions on the space of Zariski-dense leaves of $X_{\tau}$; their images generate $\mathrm{Bih}(E_{\tau})$. ∎